OpenStudy (anonymous):

Find the eigenfunctions and their associated eigenvalues of the operator [-d/dx^2]

OpenStudy (anonymous):

\[-\frac{d}{dx^2}\]

OpenStudy (anonymous):

Eigenfunctions are \[e^{\sigma x}, Cos{(\nu x + \theta)}\]

OpenStudy (anonymous):

Here the three parameters are CONSTANTS \[\sigma, \nu, \theta\] The first function's eigenvalue will be -sigma^2, the second case eigenvalue will be \[\nu^2\] B.t.w. the sin is of course a special case of the \[Cos(\nu x + \theta)\]

OpenStudy (anonymous):

@henpen

OpenStudy (anonymous):

Ahh, and one minor correction to your question: one should write of course\[\frac{ d^2 }{ dx^2 }\] and not the way u wrote.

OpenStudy (anonymous):

Thank you! You are a beautiful openstudier!

OpenStudy (anonymous):

Is there any proof that there can be no other functions that 'fit'?

OpenStudy (anonymous):

Please write it as "testimonial" I will be very much obliged to you.

OpenStudy (anonymous):

Yes - there is a proof - it involves the fact that this is a Shturm-Liouville operator and these function s form a COMPLETE BASE for ALL possible solutions. For Details consult - Shturm-Liouville theory, Second Order Equations, Hilbert Sapce BAses

OpenStudy (anonymous):

'Hilbert Space Bases' are a single thing, or is it Hilbert space and bases?

OpenStudy (anonymous):

Well try both. But As a single thing - if works, is the thing

OpenStudy (anonymous):

Which of your list holds the simplest proof?

OpenStudy (anonymous):

Shturm-Liouville theory